C1:SUS-ETMX_OLPIT_GAIN set to 1.0 OLYAW 1.0
ETMY -0.2 -0.2
ITMY 2.0 -2.0
ITMX 0.5 0.5
BS 0.4 -0.4
PRM 0.5 -0.7
SRM 1.0 1.0
Earlier today Rana and I made power spectra of ETMY_OPLEV_ERROR signals with servo on and off.
It was indicating that the servo is not doing anything. These gain values were not set since IFO rebuilt.
Valera's entries were searched also. He did not do such thing. Rana may know where it is in the elog if it happened.